IR SPECTROSCOPY AT THE CANADIAN LIGHT SOURCE: THE ν_{11} FUNDAMENTAL AND ν_{16} + ν_{18} - ν_{18} HOT BAND OF TRANS-ACROLEIN

HONG-YU SHI, \underline{LI-HONG XU}, R.M. LEES, Centre for Laser, Atomic and Molecular Sciences (CLAMS), Department of Physics, University of New Brunswick, Saint John, NB, Canada E2L 4L5; D.W. TOKARYK, CLAMS, Department of Physics, University of New Brunswick, Fredericton, NB, Canada E3B 5A3; A.R.W. McKELLAR, Steacie Institute for Molecular Sciences, National Research Council, Ottawa, ON, Canada K1A 0R6; D.R.T. APPADOO, Canadian Light Source, 101 Perimeter Road, University of Saskatchewan, Saskatoon, SK, Canada S7N 0X4.

The ν_{11} fundamental and the ν_{16} + ν_{18} - ν_{18} hot band have been identified in the high-resolution Fourier transform spectrum of trans-acrolein (CH_{2}=CH-CH=O) recorded in the 10-μm region on the Far-Infrared beamline of the Canadian Light Source synchrotron in Saskatoon. The two bands are centered at 912 cm^{−1} and 957 cm^{−1}, respectively, with the excited ν_{11} state corresponding to the A′ in-plane CH_{2}-rocking mode and the ν_{16} +ν_{18} state to the combination of the A′′ out-of-plane CH_{2} rock with the low-frequency (158 cm^{−1}) A′′ C-C torsional mode.This extends our previous work on the ν_{14} and ν_{16} fundamentals centered at 993 and 959 cm^{−1}, leaving only the ν_{15} mode now to be assigned at high resolution in the 10-μm spectral region for this important atmospheric pollutant. The ν_{11} band is type a/b and the ν_{16} +ν_{18} - ν_{18} hot band is c-type, so that assignments could be confirmed by lower-state combination differences. The assigned transitions have been fitted to a Watson asymmetric rotor Hamiltonian, and molecular parameters for both states will be reported.